OpenStudy (anonymous):

Complex number arithmetic. evaluate: e^(1 + i*pi/4) final result in cartesian form but im confused on where to start. I know its equivalent to e*e^(i*pi/4) I dont see how that would help me tho.

OpenStudy (aum):

\[\large e^{1 + i\pi/4} = e * e^{i\pi / 4} = e * (\cos\pi/4 + i\sin\pi/4) = ?\]

OpenStudy (anonymous):

= e*(cos(pi/4)) + e*i*sin(pi/4). then x = ecos(pi/4) = (e*sqrt(2))/2 y = esin(pi/4) = same as above thus, z= (e*sqrt(2))/2 + i(e*sqrt(2))/2 ??

OpenStudy (aum):

yes.

OpenStudy (anonymous):

great thanks I forgot about that essetial identity...

OpenStudy (aum):

You are welcome.